Q8) You observe a stock price of S 18.75. You expect a dividend growth rate of 5%, and the most recent dividend was $1.50. What is the required return? What is the dividend yield and capital gains yield? Do 1 +g) 0
Answer #1
A B C D E F G H I J
2
3
4 As per dividend growth model, Price of share is the present value of all future dividends discounted at cost of equity.
5 Current Price of Share (P) =Div0*(1+g)/(r-g)
6 Where Div1 is the dividend paid next year, r is cost of equity and g is perpetual growth rate of dividend.
7
8 Cost of equity can be calculated as below:
9 Cost of equity, r(E) = (Div0*(1+g)/P)+g
10
11 Given the following data:
12 Dividend (Div0) $1.50
13 Price (P) $18.75
14 Growth rate (g) 5.0%
15
16 From Dividend growth model,
17 Cost of equity, r(E) = (Div0*(1+g)/P)+g
18 Cost of equity= 13.40% =(D12*(1+D14)/D13)+D14
19
20 Hence required return is 13.40%
21
22 Dividend yield = Dividend Paid /Current Price of share
23 8.00% =D12/D13
24
25 Hence Dividend Yield is 8.00%
26
27 Capital gain yield at Year 0 =(Price at Year 1 - Price at Year 0)/Price at year 0
28
29 Calculation of Share Price at Year 1:
30
31 Given the following data:
32 Dividend (Div1) $1.58 =D12*(1+D14)
33 Cost of equity (r) 13%
34 Growth rate (g) 5.0%
35
36 From Dividend growth model,
37 Share Price at year 1, P1 =Div3*(1+g)/(r-g)
38 $19.69 =D32*(1+D34)/(D33-D34)
39
40 Hence Share Price at Year 1 is $19.69
41
42 Capital gain yield at Year 0 =(Price at Year 1 - Price at Year 0)/Price at year 0
43 5.00% =(D40-D13)/D13
44
45 Hence the capital gain yield at Year 0 is 5.00%
46